111. A realtor Selya Cribb, has been told tha 42% of homeowners in a city prefer to have a basement. he surveys a group of 250 homeowners randomly chosen from her client list. 11. Find the mean of the sampling distribution model of the proportion of homeowners in this sample uho prefer a basement. 13) 0.028 C)105 D) 0001 E) 71 12. Find the standard error of the sampling distribution of the proportion of homeowners in this sample who prefer a basement. A) 0.028 B) 117C) 0.42 0.031 E) 71 13. If only 40% of the homeowners in the survey prefer a basement (100 of the 250), should the realtor doubt the information she was originally given? Why or why not? A) Yes. The probability of getting these results is 0.142, which is unusual. B) Yes. The probability of getting these results is 0.259, which is unusual. C) No. The probability of getting these results is 0.142, which is not unusual. D) No. The probability of getting these results is 0.259, which is not unusual. E) Yes. All samples should result in 39% of homeowners who prefer a basement. IV. Choose the best response. 14. An archer is able to hit the bulls eye 65% of the time. He enters a competition where he shoots 8 arrows. Let the random variable X be the number of buls-eyes he hits during the competition What is the probability that he gets exactly 5 bulls-eyes? A) 0.279 B) 0.74 C) 5.2 D) 0.572 E) 0.145 15. At a large block party there are 290 men and 120 women. You want to ask for opinions about how to improve the next party. You decide to randomly choose 15 men and 15 women and ask thema series of questions. What is the sample being considered in the study? A) the 15 men and 15 women questioned B) All of the people at the party C) The percentage of women at the party D) the most frequently recommended improvement E) the location of the party 16. Find the z-score for the value 100, when the mean is 66 and the standard deviation is 8 A) z-4.13 B) z 4.25 C) z--1.39 D) z 1.39 E) z--34 Pa
Answer #1

Ans:

11)mean=0.42

12)standard error=sqrt(0.42*(1-0.42)/250)=0.031

13)

z=(0.4-0.42)/0.031

z=-0.645

P(z<-0.645)=0.259

Option D is correct.(No,as 0.259,so not unusual)

14)

P(x=5)=8C5*0.655*0.353=0.279

15)the 15 men and 15 women questioned.

16)z=(100-66)/8

z=34/8

z=4.25